(2 points) Calculate the pH of each of the following solutions at 25 oC: 0.250 M HCl 0.250 M NH3 0.250 M C6H5COOH (be...

  1. (2 points) Calculate the pH of each of the following solutions at 25 oC:
  1. 0.250 M HCl
  2. 0.250 M NH3
  3. 0.250 M C6H5COOH (benzoic acid)
  4. 0.250 M NH4Cl
  5. 0.250 M KOH
  6. 0.250 M CH3COOH (acetic acid ) + 0.200 M CH3CO2Na (sodium acetate)
Answer #1

06-2sъм на hice strong acid ри - los (o250) ри - о-6010-250M NUzolt py weak base Kb = 186 x 103 ki 18 at 25°C Traslog C Pkb=4.74 04 = { [ 4.74 - 1990.250) [4-74-060] a 2.07 pu =Pka=4-2 7 0.250M benzoic acid. ka = 6.3X10 (maak acid) Py a 2 [Pika - loge - [ 4-2-060] - ₂ x 3.6d) 0.250M Nuace ttce Nuqt (S) Nyce T (w) S Kb = 18810 IN MO Nyt u Au utt Pon = √ [14 - Pub - loge -2 [14-4-74 – 0.60 - 4.33 1е) 0-2 50 м Кои 100 song bоа с. Po4 - by (o 2го) = 0:00 4 – 12 401В. 250 м сч, Софи - о. 2 оо мом, Coolvo Acidic buffer Cool to go Ка? - Аб PH = pka + log [Cuzcoo] [ Сиз Cook) Henderson P4 -